OpenStudy (anonymous):

The function y = –16t^2 + 224 models the height y in feet of a stone t seconds after it is dropped from the edge of a vertical cliff. How long will it take the stone to hit the ground? Round to the nearest hundredth of a second. 0.25 seconds 5.29 seconds 7.48 seconds 3.74 seconds

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

If y is the height of the stone when it hits the ground, then what does y equal?

OpenStudy (e.cociuba):

So... Wat do u think??

OpenStudy (anonymous):

Im confused on how to solve -16t^2+224. Can you explain to me how to do it so that I can do it on the next problem lol I think I need to start there

OpenStudy (anonymous):

So it should look like 0=-16t^2 +, wait how'd you get 486?

OpenStudy (e.cociuba):

Oops I mean: Set y to zero and solve for t: y = -16t^2 + 224 MY bad...

OpenStudy (e.cociuba):

Then u do: 0 = -16t^2 + 224 16t^2 = 224 Can u solve from here?...

OpenStudy (anonymous):

would you divde 224 but -16?

OpenStudy (e.cociuba):

Not by -16.. 1) add 16t^2 to both sides 2) divide by 16 3) And then take the square root

OpenStudy (anonymous):

so then its 0=-16t^2 +224+ -16t^2/ 16

OpenStudy (e.cociuba):

0 = -16t^2 + 224 16t^2 = 224 1.) You divide by 16 on each side. So it would be 224/16= 14 2.) Then u take the square root of 14 3.) Then answer I got was 3.74165 (round and u get D. (3.74 seconds))

OpenStudy (e.cociuba):

Does tht make sence??...

OpenStudy (anonymous):

Oh yes it does I see what I was doing wrong. I just put too many -16t^2's! Thank you so much!! :))

OpenStudy (e.cociuba):

Haha Np:)